Hi. May I have some help with the following equation?
Grain is falling from a chute onto the ground, forming a conical pile whose diameter is always three times its height. How high is the pile (to the nearest hundredth of a foot) when it contains 900 ft3 of grain?

Answer by stanbon(75887)

Grain is falling from a chute onto the ground, formi9x^2ng a conical pile whose diameter is always three times its height. How high is the pile (to the nearest hundredth of a foot) when it contains 900 ft3 of grain?
-------------------------
Let the height be "x" ft. ; The diameter is "3x" : radius = (3x/2)
--------------------
Volume = (pi)r^2 h
900 = (pi)(3x/2)^2*x
900 = (pi)(9x^2/4)x
900/(pi) = (9/4)x^3
(4/9)(900/pi)= x^3
400/pi = x^3
x = 5.0308 ft (height)
=============================
Cheers,
Stan H.